# Waterloo, Ontario, N2L 2Y5, Canada

### v1

@@ Abstract
The connected wedge theorem May:2019odp ; May:2022clu states that in order to have a scattering process in the bulk, it is necessary to have $O(1/G_{N})$ mutual information between certain “decision” regions in the boundary theory. While this large mutual information is not generally sufficient to imply scattering, Caminiti:2024ctd showed that for a certain class of geometries, bulk scattering is implied by a certain relation between two (possibly non-minimal) Ryu–Takayanagi surfaces. Here, we show that the 2-to-2 version of the theorem becomes an equivalence in pure AdS3: large mutual information between appropriate boundary subregions is both necessary and sufficient for bulk scattering. This result allows us to extend the findings of Caminiti:2024ctd to a broader class of asymptotically AdS3 spacetimes, which we illustrate with the spinning conical defect geometry. In contrast, we find that matter sources can disrupt this converse relation, and that the $n$-to-$n$ version of the theorem with $n>2$ lacks a converse even in the AdS3 vacuum.
@@ 1 Introduction
The key result underlying these investigations is the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef (see also Rangamani:2016dms ), which relates the entanglement entropy of a boundary subregion to the area of a spacelike extremal surface in the bulk spacetime. Detailed calculations have shown that at least in certain settings, the entanglement structure of a boundary timeslice determines the spacetime geometry of the corresponding Cauchy surface in the bulk Balasubramanian:2013lsa ; Myers:2014jia ; Czech:2014wka ; Balasubramanian:2014sra ; Headrick:2014eia . Further, making use of a ‘first law’ of entanglement entropy Blanco:2013joa , it can be shown Lashkari:2013koa ; Faulkner:2013ica ; Swingle:2014uza ; Faulkner:2017tkh that the Ryu-Takayanagi prescription requires that spacetimes dual to small excitations of the CFT vacuum state must satisfy Einstein’s equations perturbatively expanded around pure AdS spacetime. Hence holographic entanglement entropy plays a key role in understanding both the bulk geometry and its dynamics.
The important role of entanglement in the dynamics and causal structure of the bulk theory was highlighted in recent work on holographic scattering May:2019yxi ; May:2019odp ; May:2021nrl ; May:2022clu . The essential result May:2019odp ; May:2022clu is summarized as the connected wedge theorem (CWT), which observes that strong entanglement of two or more boundary subregions is a necessary precondition for certain bulk processes involving timelike-separated input and output regions. The CWT is proven for general asymptotically AdS3 geometries where the HRRT formula Hubeny:2007xt evaluates the entanglement entropy in the boundary theory and the bulk obeys the null energy condition. [^1] The CWT also extends to semiclassical spacetimes with quantum matter May:2019odp ; May:2022clu using generalized entropy and quantum extremal surfaces Faulkner:2013ana ; Engelhardt:2014gca , and assuming the quantum focusing conjecture Bousso:2015mna and quantum maximin formula Akers:2019lzs .
@@
Figure 1: (a) By the CWT, a bulk-only scattering process from $c_{1},c_{2}$ to $r_{1},r_{2}$ requires associated boundary regions (green) to be strongly correlated, as captured by the connectivity of the entanglement wedge. (b) When there is no bulk-only scattering, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ need not be strongly correlated. In particular, they can have disconnected minimal-area extremal surfaces. This figure is reproduced from May:2019yxi .
Hence, from the boundary perspective, the process is enabled by the strong entanglement of the degrees of freedom in the two decision regions. The intuition behind this result is the following: The process which proceeds by a local interaction of the probes in in the bulk must somehow be encoded in the boundary CFT by the basic tenets of holography. However, since the probes cannot actually meet in the boundary theory, this process must be encoded in a nonlocal manner. In the context of quantum information, it has been understood that nonlocal protocols to achieve certain quantum computation tasks require specific entanglement resources – see e.g., Buhrman:2014aou ; Kent:2011jim . Hence, one is led to expect that bulk scattering processes can be encoded as a nonlocal quantum protocol in the boundary, requiring some minimum amount of entanglement among particular boundary subregions as a resource. The connected entanglement wedge emerging from the CWT implies the required amount of entanglement is characterized by a mutual information $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ of order ${\cal O}(1/G_{N})$.
We emphasize that the CWT does not provide an if-and-only-if statement. That is, while large mutual information (i.e., a connected wedge) is necessary for these bulk-only scattering processes, it is not sufficient. Counterexamples are known May:2019odp ; Caminiti:2024ctd , yielding the intuition that not all correlations contributing to $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ are useful in facilitating the nonlocal quantum protocols. Further, counterexamples to the converse also exist for a stronger version of the CWT known as the regions theorem May:2021nrl , which will be introduced in section 2 and is our main focus in this work. The absence of a converse raises the question: what boundary quantity, if any, serves as a diagnostic for bulk scattering? One hopes that answering this question will improve our understanding of the structure of entanglement needed for the corresponding nonlocal quantum protocols.
@@
Figure 2: The 2-to-2 scattering setup on the boundary cylinder, which was considered in Caminiti:2024ctd . Both input regions, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, are on a fixed time slice, have the same size $x$, and their midpoints are separated by an angle $\theta$. The output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$, are maximized to cover an entire Cauchy slice. Note that the vertical edges on the left and and right are identified.
@@
Towards this goal, Caminiti:2024ctd examined the regions version of the CWT in asymptotically AdS3 spacetimes containing a (static) massive object, namely a conical defect or a BTZ black hole. Figure 2 illustrates the 2-to-2 scattering setup on the boundary cylinder considered in Caminiti:2024ctd . The two input regions $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ lie on a fixed time slice and have the same size $x$, and their midpoints are separated by an angle $\theta$. The output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$, are maximized to cover an entire Cauchy slice, as illustrated. In this situation, there are three possible topologies for the entanglement wedge ${\cal E}({\hat{\mathcal{V}}_{1}\cup\hat{\mathcal{V}}_{2}})$: a disconnected topology (denoted $d$) and two connected topologies (denoted $u$ and $o$). The $u$ and $o$ candidates differ by whether or not the entanglement wedge does ($o$) or does not ($u$) include the massive object at the center of the spacetime – see figure 3.
@@
Figure 3: A constant-$t$ slice of the conical defect geometry showing, from left to right, the $d$, $u$, and $o$ candidates for $\mathcal{E}(A\cup B)$. Figure copied from Caminiti:2024ctd .
@@
The analysis in Caminiti:2024ctd revealed that a curious relationship [^2] The inequality $u<d$ is a convenient shorthand to indicate that the area of the extremal surface associated with the $u$ configuration is smaller than that of the corresponding $d$ configuration. $u<d$ was sufficient to imply holographic scattering in these nontrivial AdS3 backgrounds. Of particular note was that the $u<d$ condition still implied scattering even when $u$ did not describe the actual entanglement wedge, i.e., the corresponding extremal surface was not the minimal RT surface. In the present work, we test the robustness of the $u<d$ converse to the CWT with a view towards understanding better what kind of entanglement allows for holographic scattering.
@@
To summarize our results, we begin by observing that in pure AdS3, the 2-to-2 CWT has a converse, as anticipated in May:2021nrl . We then explore the consequences of this result for holographic scattering in a general asymptotically AdS3 vacuum solution, $\mathcal{M}$. Our key insight is that any open, simply connected subregion $\mathcal{N}\subset\mathcal{M}$ which does not contain defects or singularities is equivalent to a portion of pure AdS3, and so the CWT converse in pure AdS3 has implications for this broader class of geometries. Building on this insight, we generalize the $u<d$ result of Caminiti:2024ctd to all asymptotically AdS3 vacuum solutions. Specifically, we show that a $u<d$-type relation among nonminimal surfaces contained in $\mathcal{N}$ implies holographic scattering in pure states. Notably, this extension applies to, e.g., spinning defect or spinning BTZ backgrounds, with arbitrary choices of spacelike separated input regions (not only those shown in figure 2).
@@
Lastly, we extend our analysis to the $n$-to-$n$ CWT May:2022clu . We find that for $n>2$, the CWT does not admit a converse even in the AdS3 vacuum. This indicates that $u<d$-type inequalities should not be expected to govern higher-$n$ scattering.
The remainder of the paper is organized as follows: In section 2, we review the connected wedge theorem and demonstrate that the 2-to-2 CWT naturally has a converse in pure AdS3. Building on this observation, in section 3, we generalize the $u<d$ result of Caminiti:2024ctd to general asymptotically AdS3 vacuum solutions. In section 4, we show that the $u<d$ result fails to generalize for asymptotically AdS3 solutions with matter by analyzing holographic scattering in a simple background containing a spherical shell. In section 5, we extend our analysis to consider the $n$-to-$n$ CWT and demonstrate that there is no converse with $n>2$, even in pure AdS3. Finally, we discuss our results and possible future directions in section 6, in particular drawing connections to future work ap ; ap2 . Additional technical details are provided in the appendices: Appendix A provides useful expressions for spacelike and null geodesics in global AdS3 coordinates. Appendix B examines how the results of section 3 apply for the spinning conical defect. Appendix C contains further details on the shell geometry and the calculations discussed in section 4.
@@ 2 The connected wedge theorem in pure AdS3
In this section, we introduce the regions-based connected wedge theorem May:2021nrl and show that for 2-to-2 scattering in pure AdS3, the converse of the connected wedge theorem always holds: namely, having a connected entanglement wedge is enough to imply holographic scattering is possible.
@@ 2.1 The connected wedge theorem
We state the connected wedge theorem (CWT) May:2021nrl :
@@
Assume that ${\hat{\mathcal{C}}}_{i}\subseteq\hat{\mathcal{V}}_{i}$. Assume also that the bulk geometry obeys the null energy condition and the HRRT surface can be found using the maximin formula Wall:2012uf . Define the bulk scattering region
@@
Note that the example discussed in the introduction and depicted in figure 1 involves a special case of the CWT where the boundary regions consist of single points: $\hat{\mathcal{C}}_{1}=\{c_{1}\}$, $\hat{\mathcal{R}}_{2}=\{r_{2}\}$, etc. This points version of the theorem was developed first May:2019odp , but we focus on the regions theorem throughout the main text because it is more general. We henceforth reserve the labels $c_{i}$ to refer to the past-most points of the $\hat{\mathcal{C}}_{i}$ regions, and $r_{i}$ to refer to the future-most points of the $\hat{\mathcal{R}}_{i}$ regions. We emphasize that the two versions of the CWT are the same in pure AdS3, the case of interest in this section, because the equivalence of causal wedges and entanglement wedges in pure AdS3 implies
@@ 2.2 The converse in pure AdS: 2-to-2 case
We now demonstrate that in pure AdS3, holographic scattering is not only necessary but also sufficient for a connected entanglement wedge. This was demonstrated in appendix B of May:2021nrl for the case where $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ are the causal developments of equally-sized intervals on a constant time-slice; our approach is very similar to the approach in this reference. [^3] We thank Alex May for conversations on this topic.
@@
We would like to compare eq. (15) to the condition for a non-empty bulk scattering region, i.e., $J_{1,2\to 1,2}\neq\emptyset$. We can simplify the expression (2) for the scattering region by observing that $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ share the same RT surface: $\gamma_{\hat{\mathcal{R}}_{1}}=\gamma_{\hat{\mathcal{R}}_{2}}$. Hence, [^6] For a more detailed justification of eq. (16), see Caminiti:2024ctd .
@@
Here $c_{1}$ and $c_{2}$ correspond to the past-most points of $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, respectively – see figure 4. Now, eq. (17) says that holographic scattering is possible if and only if two probe systems leaving from $c_{1}$ and $c_{2}$ respectively can meet at $\gamma_{\hat{\mathcal{R}}_{1}}$. The threshold for scattering thus occurs when null rays emanating from $c_{1}$ and $c_{2}$ intersect precisely at $\gamma_{\hat{\mathcal{R}}_{1}}$. Note that $\gamma_{\hat{\mathcal{R}}_{1}}$ is contained in the future lightsheet emanating from the past-most point of $\hat{\mathcal{R}}_{1}$, which we have positioned at the origin $(U,V,z)=(0,0,0)$. Hence, we can determine the scattering threshold by considering three lightcones emanating from the conformal boundary at $z=0$, with vertices $(0,0)$, $c_{1}=(U_{1},V_{1})$, and $c_{2}=(U_{1},V_{1})$ respectively. These lightcones are given by

### v3

@@ Abstract
The connected wedge theorem May et al. (2020, 2022) states that in order to have a scattering process in the bulk, it is necessary to have $O(1/G_{N})$ mutual information between certain “decision” regions in the boundary theory. While this large mutual information is not generally sufficient to imply scattering, Caminiti et al. (2024) showed that for a certain class of geometries, bulk scattering is implied by a certain relation between two (possibly non-minimal) Ryu–Takayanagi surfaces. Here, we show that the 2-to-2 version of the theorem becomes an equivalence in pure AdS3: large mutual information between appropriate boundary subregions is both necessary and sufficient for bulk scattering. This result allows us to extend the findings of Caminiti et al. (2024) to a broader class of asymptotically AdS3 spacetimes, which we illustrate with the spinning conical defect geometry. In contrast, we find that matter sources can disrupt this converse relation, and that the $n$-to-$n$ version of the theorem with $n>2$ lacks a converse even in the AdS3 vacuum.
@@ 1 Introduction
The key result underlying these investigations is the Ryu-Takayanagi (RT) formula Ryu and Takayanagi (2006b, a) (see also Rangamani and Takayanagi (2017)), which relates the entanglement entropy of a boundary subregion to the area of a spacelike extremal surface in the bulk spacetime. Detailed calculations have shown that at least in certain settings, the entanglement structure of a boundary timeslice determines the spacetime geometry of the corresponding Cauchy surface in the bulk Balasubramanian et al. (2014a); Myers et al. (2014); Czech et al. (2014); Balasubramanian et al. (2015); Headrick et al. (2014). Further, making use of a ‘first law’ of entanglement entropy Blanco et al. (2013), it can be shown Lashkari et al. (2014); Faulkner et al. (2014); Swingle and Van Raamsdonk (2014); Faulkner et al. (2017) that the Ryu-Takayanagi prescription requires that spacetimes dual to small excitations of the CFT vacuum state must satisfy Einstein’s equations perturbatively expanded around pure AdS spacetime. Hence holographic entanglement entropy plays a key role in understanding both the bulk geometry and its dynamics.
The important role of entanglement in the dynamics and causal structure of the bulk theory was highlighted in recent work on holographic scattering May (2019); May et al. (2020); May (2021); May et al. (2022). The essential result May et al. (2020, 2022) is summarized as the connected wedge theorem (CWT), which observes that strong entanglement of two or more boundary subregions is a necessary precondition for certain bulk processes involving timelike-separated input and output regions. The CWT is proven for general asymptotically AdS3 geometries where the HRRT formula Hubeny et al. (2007) evaluates the entanglement entropy in the boundary theory and the bulk obeys the null energy condition. [^1] The CWT also extends to semiclassical spacetimes with quantum matter May et al. (2020, 2022) using generalized entropy and quantum extremal surfaces Faulkner et al. (2013); Engelhardt and Wall (2015), and assuming the quantum focusing conjecture Bousso et al. (2016) and quantum maximin formula Akers et al. (2020).
@@
Figure 1: (a) By the CWT, a bulk-only scattering process from $c_{1},c_{2}$ to $r_{1},r_{2}$ requires associated boundary regions (green) to be strongly correlated, as captured by the connectivity of the entanglement wedge. (b) When there is no bulk-only scattering, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ need not be strongly correlated. In particular, they can have disconnected minimal-area extremal surfaces. This figure is reproduced from May (2019).
Hence, from the boundary perspective, the process is enabled by the strong entanglement of the degrees of freedom in the two decision regions. The intuition behind this result is the following: The process which proceeds by a local interaction of the probes in in the bulk must somehow be encoded in the boundary CFT by the basic tenets of holography. However, since the probes cannot actually meet in the boundary theory, this process must be encoded in a nonlocal manner. In the context of quantum information, it has been understood that nonlocal protocols to achieve certain quantum computation tasks require specific entanglement resources – see e.g., Buhrman et al. (2014); Kent et al. (2011). Hence, one is led to expect that bulk scattering processes can be encoded as a nonlocal quantum protocol in the boundary, requiring some minimum amount of entanglement among particular boundary subregions as a resource. The connected entanglement wedge emerging from the CWT implies the required amount of entanglement is characterized by a mutual information $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ of order ${\cal O}(1/G_{N})$.
We emphasize that the CWT does not provide an if-and-only-if statement. That is, while large mutual information (i.e., a connected wedge) is necessary for these bulk-only scattering processes, it is not sufficient. Counterexamples are known May et al. (2020); Caminiti et al. (2024), yielding the intuition that not all correlations contributing to $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ are useful in facilitating the nonlocal quantum protocols. Further, counterexamples to the converse also exist for a stronger version of the CWT known as the regions theorem May (2021), which will be introduced in section 2 and is our main focus in this work. The absence of a converse raises the question: what boundary quantity, if any, serves as a diagnostic for bulk scattering? One hopes that answering this question will improve our understanding of the structure of entanglement needed for the corresponding nonlocal quantum protocols.
@@
Figure 2: The 2-to-2 scattering setup on the boundary cylinder, which was considered in Caminiti et al. (2024). Both input regions, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, are on a fixed time slice, have the same size $x$, and their midpoints are separated by an angle $\theta$. The output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$, are maximized to cover an entire Cauchy slice. Note that the vertical edges on the left and and right are identified.
@@
Towards this goal, Caminiti et al. (2024) examined the regions version of the CWT in asymptotically AdS3 spacetimes containing a (static) massive object, namely a conical defect or a BTZ black hole. Figure 2 illustrates the 2-to-2 scattering setup on the boundary cylinder considered in Caminiti et al. (2024). The two input regions $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ lie on a fixed time slice and have the same size $x$, and their midpoints are separated by an angle $\theta$. The output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$, are maximized to cover an entire Cauchy slice, as illustrated. In this situation, there are three possible topologies for the entanglement wedge ${\cal E}({\hat{\mathcal{V}}_{1}\cup\hat{\mathcal{V}}_{2}})$: a disconnected topology (denoted $d$) and two connected topologies (denoted $u$ and $o$). The $u$ and $o$ candidates differ by whether or not the entanglement wedge does ($o$) or does not ($u$) include the massive object at the center of the spacetime – see figure 3.
@@
Figure 3: A constant-$t$ slice of the conical defect geometry showing, from left to right, the $d$, $u$, and $o$ candidates for $\mathcal{E}(A\cup B)$. Figure copied from Caminiti et al. (2024).
@@
The analysis in Caminiti et al. (2024) revealed that a curious relationship [^2] The inequality $u<d$ is a convenient shorthand to indicate that the area of the extremal surface associated with the $u$ configuration is smaller than that of the corresponding $d$ configuration. $u<d$ was sufficient to imply holographic scattering in these nontrivial AdS3 backgrounds. Of particular note was that the $u<d$ condition still implied scattering even when $u$ did not describe the actual entanglement wedge, i.e., the corresponding extremal surface was not the minimal RT surface. In the present work, we test the robustness of the $u<d$ converse to the CWT with a view towards understanding better what kind of entanglement allows for holographic scattering.
@@
To summarize our results, we begin by observing that in pure AdS3, the 2-to-2 CWT has a converse, as anticipated in May (2021). We then explore the consequences of this result for holographic scattering in a general asymptotically AdS3 vacuum solution, $\mathcal{M}$. Our key insight is that any open, simply connected subregion $\mathcal{N}\subset\mathcal{M}$ which does not contain defects or singularities is equivalent to a portion of pure AdS3, and so the CWT converse in pure AdS3 has implications for this broader class of geometries. Building on this insight, we generalize the $u<d$ result of Caminiti et al. (2024) to all asymptotically AdS3 vacuum solutions. Specifically, we show that a $u<d$-type relation among nonminimal surfaces contained in $\mathcal{N}$ implies holographic scattering in pure states. Notably, this extension applies to, e.g., spinning defect or spinning BTZ backgrounds, with arbitrary choices of spacelike separated input regions (not only those shown in figure 2).
@@
Lastly, we extend our analysis to the $n$-to-$n$ CWT May et al. (2022). We find that for $n>2$, the CWT does not admit a converse even in the AdS3 vacuum. This indicates that $u<d$-type inequalities should not be expected to govern higher-$n$ scattering.
The remainder of the paper is organized as follows: In section 2, we review the connected wedge theorem and demonstrate that the 2-to-2 CWT naturally has a converse in pure AdS3. Building on this observation, in section 3, we generalize the $u<d$ result of Caminiti et al. (2024) to general asymptotically AdS3 vacuum solutions. In section 4, we show that the $u<d$ result fails to generalize for asymptotically AdS3 solutions with matter by analyzing holographic scattering in a simple background containing a spherical shell. In section 5, we extend our analysis to consider the $n$-to-$n$ CWT and demonstrate that there is no converse with $n>2$, even in pure AdS3. Finally, we discuss our results and possible future directions in section 6, in particular drawing connections to future work Lima et al. (2025b, a). Additional technical details are provided in the appendices: Appendix A provides useful expressions for spacelike and null geodesics in global AdS3 coordinates. Appendix B examines how the results of section 3 apply for the spinning conical defect. Appendix C contains further details on the shell geometry and the calculations discussed in section 4.
@@ 2 The connected wedge theorem in pure AdS3
In this section, we introduce the regions-based connected wedge theorem May (2021) and show that for 2-to-2 scattering in pure AdS3, the converse of the connected wedge theorem always holds: namely, having a connected entanglement wedge is enough to imply holographic scattering is possible.
@@ 2.1 The connected wedge theorem
We state the connected wedge theorem (CWT) May (2021):
@@
Assume that ${\hat{\mathcal{C}}}_{i}\subseteq\hat{\mathcal{V}}_{i}$. Assume also that the bulk geometry obeys the null energy condition and the HRRT surface can be found using the maximin formula Wall (2014). Define the bulk scattering region
@@
Note that the example discussed in the introduction and depicted in figure 1 involves a special case of the CWT where the boundary regions consist of single points: $\hat{\mathcal{C}}_{1}=\{c_{1}\}$, $\hat{\mathcal{R}}_{2}=\{r_{2}\}$, etc. This points version of the theorem was developed first May et al. (2020), but we focus on the regions theorem throughout the main text because it is more general. We henceforth reserve the labels $c_{i}$ to refer to the past-most points of the $\hat{\mathcal{C}}_{i}$ regions, and $r_{i}$ to refer to the future-most points of the $\hat{\mathcal{R}}_{i}$ regions. We emphasize that the two versions of the CWT are the same in pure AdS3, the case of interest in this section, because the equivalence of causal wedges and entanglement wedges in pure AdS3 implies
@@ 2.2 The converse in pure AdS: 2-to-2 case
We now demonstrate that in pure AdS3, holographic scattering is not only necessary but also sufficient for a connected entanglement wedge. This was demonstrated in appendix B of May (2021) for the case where $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ are the causal developments of equally-sized intervals on a constant time-slice; our approach is very similar to the approach in this reference. [^3] We thank Alex May for conversations on this topic.
@@
We would like to compare eq. (15) to the condition for a non-empty bulk scattering region, i.e., $J_{1,2\to 1,2}\neq\emptyset$. We can simplify the expression (2) for the scattering region by observing that $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ share the same RT surface: $\gamma_{\hat{\mathcal{R}}_{1}}=\gamma_{\hat{\mathcal{R}}_{2}}$. Hence, [^6] For a more detailed justification of eq. (16), see Caminiti et al. (2024).
@@
Here $c_{1}$ and $c_{2}$ correspond to the past-most points of $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, respectively – see figure 4. Now, eq. (17) says that holographic scattering is possible if and only if two probe systems leaving from $c_{1}$ and $c_{2}$ respectively can meet at $\gamma_{\hat{\mathcal{R}}_{1}}$. The threshold for scattering thus occurs when null rays emanating from $c_{1}$ and $c_{2}$ intersect precisely at $\gamma_{\hat{\mathcal{R}}_{1}}$. Note that $\gamma_{\hat{\mathcal{R}}_{1}}$ is contained in the future lightsheet emanating from the past-most point of $\hat{\mathcal{R}}_{1}$, which we have positioned at the origin $(U,V,z)=(0,0,0)$. Hence, we can determine the scattering threshold by considering three lightcones emanating from the conformal boundary at $z=0$, with vertices $(0,0)$, $c_{1}=(U_{1},V_{1})$, and $c_{2}=(U_{2},V_{2})$ respectively. These lightcones are given by